### v3

# Integral points and volume of integral-integral affine manifolds

Abstract.
We give an elementary proof that, for a closed manifold with an integral-integral affine structure, its total volume and number of integral points coincide. The proof uses rational Ehrhart theory and elementary Fourier analysis to estimate the difference between the total volume and the number of integral points.
Key words and phrases:
Integral affine structure, Ehrhart
2020 Mathematics Subject Classification:
Primary 53C15; Secondary 52B20, 53D50
1. Introduction
On an $n$-dimensional manifold $M$, an atlas whose transition maps locally have the form
(1.1) $x\mapsto Ax+b\quad\text{ with $A\in\mathsf{GL}(n,\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$ }$
defines an integral-integral affine structure [undefi] (also called a strongly integral affine structure [undefo]). This structure determines a measure on $M$ and a subset $M_{\mathbb{Z}}\subset M$ of integral points. (See Definition 2.5, Remark 2.6, Remark 2.7, and Definition 2.12.)
For example, if $\Gamma$ is group of transformations of $\mathbb{R}^{n}$ of the form (1.1) whose action on $\mathbb{R}^{n}$ is free and proper then the quotient $M:=\mathbb{R}^{n}/\Gamma$ inherits an integral-integral affine structure. We do not know if every compact integral-integral affine manifold must be isomorphic to such a quotient; this is a special case of the Markus conjecture [undefl] (see the discussion in [undefe]).
In this paper we give an elementary proof of the following theorem:
1.2 Theorem.
Let $M$ be a compact integral-integral affine manifold. Then its number of integral points is equal to its total volume.
Our motivation for this project comes from geometric quantization. Roughly, geometric quantization provides recipes to pass from a symplectic manifold $(X,\omega)$, which serves as a model for phase space in classical mechanics, to a Hilbert space $\mathcal{H}$, whose elements we view as wave functions in the corresponding quantum mechanical system. Such recipes involve the choice of a polarization on $X$. Properties of the quantum mechanical system often turn out to be independent of this choice; this phenomenon is called “independence of polarization”.
One example of a polarization is a compatible Kähler structure on $X$; some recipes allow a compatible almost complex structure; the dimension of the corresponding Hilbert space is typically given by the Riemann-Roch number. Another example of a polarization is a Lagrangian fibration $X\to B$. In the presence of prequantization, the base $B$ then acquires an integral-integral affine structure, and the dimension of the corresponding Hilbert space that arises from this polarization is given by the number of integral points. The Riemann-Roch number of $(X,\omega)$ then coincides with the volume of $B$; thus, Theorem 1.2 becomes an instance of “independence of polarization”.
Theorem 1.2 appeared as the “Downstairs Theorem” in [undefi]. Its proof was surprizingly nontrivial, and involved intersection theory on a bundle of tori over $M$. (Warning: in [undefi], the integral-integral manifold was called $B$; the letter $M$ was used for various types of torus bundles over $B$.) The proof in [undefi] was inspired by ideas that were sketched inside papers on geometric quantization and on mirror symmetry ([undef, Proof of Thm. 4.1]; [undeff, Conjecture 4.9]); the details turned out to require a subtle cohomological argument, which was provided in [undefi]. In light of this earlier work, the elementary proof that we provide in this current paper is particularly satisfying.
In this paper we provide a more elementary proof of Theorem 1.2 that relies on rational Ehrhart theory in $\mathbb{R}^{n}$ [undefd, undefa]. Here is a sketch of our argument. We introduce the lattice count function, which associates to each $m\in\mathbb{N}$ the number $L_{M}(m)$ of points in $M$ whose coordinates are in $m^{-1}\mathbb{Z}$. For a rational convex polytope $P$ in $\mathbb{R}^{n}$, the number of points in $P\cap m^{-1}\mathbb{Z}^{n}$ is quasi-polynomial in $m$ (Definition 2.10), by rational Ehrhart theory. For a compact integral-integral affine manifold $M$, we do not know how to decompose $M$ into rational convex polytopes; however, we can cover $M$ by overlapping rational convex polytopes, and apply an inclusion-exclusion argument, to conclude that $L_{M}(m)$ is a quasi-polynomial in $m$. On the other hand, by “decomposing” $M$ using a smooth partition of unity supported in charts, we conclude that $L_{M}(m)-\text{vol}(M)m^{n}=o(1)$ as $m\to\infty$. From these two results, we conclude that $L_{M}(m)=\text{vol}(M)m^{n}$ on the nose, yielding Theorem 1.2.
Acknowledgement
We are grateful to Roei Raveh for providing the idea of the proof of Proposition 3.7.
Y. Karshon’s research is partly funded by the Natural Sciences and Engineering Research Council of Canada Discovery RGPIN-2024-05798 and by the United States – Israel Binational Science Foundation Grant 2021730. O. Elisha is partly supported by Yaron Ostrover’s ISF Grant 938/22.
2. Preliminaries
2.1 Definition.
Let $M$ be a (smooth) $n$ dimensional manifold. An open cover $\mathcal{U}\coloneq\{U_{i}\}_{i\in I}$ of $M$ is called a good cover if every nonempty finite intersection $\bigcap_{j=0}^{m}{U_{j}}$ of sets $U_{j}$ from $\mathcal{U}$ is diffeomorphic to $\mathbb{R}^{n}$.
2.2 Notation.
We denote by $\textbf{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$ the set of maps $\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $x\mapsto Ax+b$ with $A\in\mathsf{GL}(n,\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$.
2.3 Definition.
A map $f\colon\Omega\rightarrow\mathbb{R}^{n}$ on an open subset $\Omega\subseteq\mathbb{R}^{n}$ is called locally integral-integral affine if each point has a neighbourhood on which the map agrees with an element of $\operatorname{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$.
2.4 Proposition.
Let $f\colon\Omega\rightarrow\mathbb{R}^{n}$ be a locally integral-integral affine map on an open subset $\Omega$ of $\mathbb{R}^{n}$. If $\Omega$ is connected, then $f$ is the restriction to $\Omega$ of some element of $\operatorname{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$.
Proof.
Write $f=(f_{1},\ldots,f_{n})$. The partial derivatives $\partial_{i}f_{j}$ are locally constant maps on the connected open set $\Omega$, so they are constant. Let $A\in\mathsf{GL}_{n}(\mathbb{Z})$ be the matrix of partial derivatives. Then $x\mapsto f(x)-Ax$ is a function with vanishing partial derivatives on the connected open set $\Omega$, so it is constant. ∎
2.5 Definition.
Let $M$ be an $n$ dimensional manifold. An atlas
$\{\varphi_{i}\colon U_{i}\rightarrow\Omega_{i}\subseteq\mathbb{R}^{n}\}_{i\in I}$
for M is called an integral-integral affine atlas if $\varphi_{j}\circ\varphi_{i}^{-1}$ is locally integral-integral affine on $\varphi_{i}(U_{i}\cap U_{j})$ for every $i,j\in I$. The integral-integral affine atlases are ordered by inclusion. An integral-integral affine structure on a manifold $M$ is a maximal integral-integral affine atlas. An integral-integral affine manifold is a manifold equipped with an integral-integral affine structure. An integral chart is an element of the maximal integral-integral affine atlas of $M$.
2.6 Remark.
Every integral-integral affine atlas is contained in a unique maximal integral-integral affine atlas. This follows from the fact that locally integral-integral affine homeomorphisms between open subsets of $\mathbb{R}^{n}$ form a pseudogroup.
2.7 Remark.
Let $M$ be an $n$ dimensional integral-integral affine manifold. Since locally integral-integral affine maps preserve the standard Lebesgue measure on $\mathbb{R}^{n}$, there exists a unique (Borel-)measure on $M$ such that each integral chart is measure preserving. Integration on an integral-integral affine manifold is then done with respect to this measure.
2.8 Definition.
A set $P\subseteq\mathbb{R}^{n}$ is called a rational polytope if there exist $v_{1},\dots,v_{k}\in\mathbb{Q}^{n}$, such that $P=\operatorname{conv}\{v_{1},\dots,v_{k}\}=\{\sum_{i=1}^{k}{\alpha_{i}v_{i}}\mid\alpha_{i}\geq 0,\;\sum_{i=1}^{k}{\alpha_{i}}=1\}$.
A point $x\in P$ is called a vertex of $P$ if $x$ is an extreme point in P. That is, if $x=\lambda x_{1}+(1-\lambda)x_{2}$ for $0\leq\lambda\leq 1$ where $x_{1},x_{2}\in P$ and $x_{1}\neq x_{2}$, then $\lambda=0$ or $\lambda=1$.
2.9 Remark.
If $P=\operatorname{conv}\{v_{1},\dots,v_{k}\}$, then the set of vertices of $P$ is a subset of $\{v_{1},\dots,v_{k}\}$.
The set of natural numbers is
$\mathbb{N}\coloneq\{1,2,\ldots\}.$
2.10 Definition.
A function $f\colon\mathbb{N}\rightarrow\mathbb{R}$ is quasi-polynomial if there exist $n\in\mathbb{N}\cup\{0\}$ and functions $\{a_{k}\colon\mathbb{N}\rightarrow\mathbb{R}\}_{k=0}^{n}$ such that
  • Each $a_{k}$ is periodic with integer period, and
  • $f(m)=\sum_{k=0}^{n}{a_{k}(m)m^{k}}$.
2.11 Definition.
For a rational polytope $P\subseteq\mathbb{R}^{n}$, we define its lattice count function by
$L_{P}\colon\mathbb{N}\rightarrow\mathbb{N}\cup\{0\}$
$L_{P}(m)\coloneq\#((m^{-1}\mathbb{Z}^{n})\cap P)=|\{x\in P\mid mx\in\mathbb{Z}^{n}\}|\,.$
2.12 Definition.
Let $M$ be an $n$-dimensional integral-integral affine manifold. Its set of integral points is
[…]
Moreover, for $m\in\mathbb{N}$, denote
[…]
2.13 Lemma.
For an $n$-dimensional integral-integral affine manifold $M$, and for $m\in\mathbb{N}$,
[…]
Proof.
Let $x\in M_{m^{-1}\mathbb{Z}}$. So there exists an integral chart $\psi\colon U^{\prime}\rightarrow\Omega^{\prime}\subseteq\mathbb{R}^{n}$ whose domain contains $x$ and such that $\psi(x)\in m^{-1}\mathbb{Z}^{n}$. Let $\phi\colon U\to\Omega\subseteq\mathbb{R}^{n}$ be another integral chart whose domain contains $x$. By compatibility, there exists a neighbourhood of $x$ on which the composition $\phi\circ\psi^{-1}$ has the form $y\mapsto b+Ay$ for $b\in\mathbb{Z}^{n}$ and $A\in\mathsf{GL}_{n}(\mathbb{Z})$. Thus $m\phi(x)=m\phi(\psi^{-1}(\psi(x)))=m(b+A(\psi(x)))=mb+A(m\psi(x))\in\mathbb{Z}^{n}$. ∎
2.14 Definition.
Let $M$ be an integral-integral affine manifold. Its lattice count function is
$L_{M}\colon\mathbb{N}\rightarrow\mathbb{N}\cup\{0,\infty\}$
$L_{M}(m)\coloneq\#M_{m^{-1}\mathbb{Z}}\,.$
2.15 Proposition.
Let $P\subseteq\mathbb{R}^{n}$ be a rational polytope, and let $f\in\operatorname{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$. Then $f(P)$ is a rational polytope.
Proof.
Let $P=\operatorname{conv}\{q_{1},\dots,q_{k}\}$ for $q_{i}\in\mathbb{Q}^{n}$. Let $A\in\mathsf{GL}_{n}(\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$ be such that $f(x)=b+Ax$. Since $f$ is an affine map, $f(P)=\operatorname{conv}\{f(q_{1}),\dots,f(q_{k})\}=\operatorname{conv}\{b+Aq_{1},\dots,b+Aq_{k}\}$.  Because $A\in\mathsf{GL}_{n}(\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$, we have $b+Aq_{i}\in\mathbb{Q}^{n}$ for all $1\leq i\leq k$. Thus, $f(P)$ is a rational polytope. ∎
2.16 Definition.
Let $P\subset M$ be a subset of an $n$-dimensional integral-integral affine manifold $M$. Then $P$ is called a rational polytope if there exists an integral chart $\phi$ whose domain contains $P$ and such that $\phi(P)$ is a rational polytope in $\mathbb{R}^{n}$.
2.17 Proposition.
Let M be an integral-integral affine manifold, $P\subseteq M$ a rational polytope and $\psi\colon W\rightarrow\Omega\subseteq\mathbb{R}^{n}$ any integral chart such that $P\subseteq W$. Then $\psi(P)\subseteq\mathbb{R}^{n}$ is a rational polytope.
Proof.
By definition, there is an integral chart $\phi\colon U\rightarrow\Omega^{\prime}\subseteq\mathbb{R}^{n}$ such that $P\subseteq U$ and $\phi(P)$ is a rational polytope. In particular, $\phi(P)$ is connected. Let $\mathcal{O}$} be the connected component of $\phi(U\cap W)$ that contains $\phi(P)$. By Proposition 2.4, there exists $f\in\operatorname{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$ that coincides on $\mathcal{O}$} with $\psi\circ\phi^{-1}$. By Proposition 2.15, $\psi(P)=f(\phi(P))$ is a rational polytope. ∎
2.18 Proposition.
Let $P_{1},P_{2}\subseteq\mathbb{R}^{n}$ be rational polytopes. Then $P_{1}\cap P_{2}$ is either empty or a rational polytope.
Proof.
The definition of a polytope in [undefg] is equivalent to our Definition 2.8, by [undefg, Theorem 1.26]. The result is then in [undefg, Proposition 1.69]. ∎
3. Main Theorems
3.1 Proposition.
Let $M$ be a closed integral-integral affine manifold. Then its lattice count function (Definition 2.11) is finite: $L_{M}(m)<\infty$ for every $m\in\mathbb{N}$.
Proof.
Assume not; let $(x_{i})_{i=0}^{\infty}$ be a sequence of distinct points in $M_{m^{-1}\mathbb{Z}}$. Since M is compact, this sequence has a converging subsequence $(x_{i_{j}})_{j=0}^{\infty}\rightarrow x^{*}$. Let $\phi:U\rightarrow\Omega\subseteq\mathbb{R}^{n}$ be an integral chart such that $x^{*}\in U$. Then for some $N>0$ we have $x_{i_{j}}\in U$ for all $j\geq N$. Then $\phi(x_{i_{j}})$ is a convergent sequence of distinct elements of $m^{-1}\mathbb{Z}^{n}$}, which is a contradiction. ∎
3.2 Proposition (Rational Ehrhart).
Let $P\subseteq\mathbb{R}^{n}$ be a rational polytope. Then its lattice count function is quasi-polynomial.
Proof.
[undefa, Proposition 3.23]. ∎
3.3 Lemma.
Let $M$ be an integral-integral affine manifold and $P_{1},P_{2}\subseteq M$ rational polytopes, and let $\phi_{1}:U_{1}\rightarrow\Omega_{1}\subseteq\mathbb{R}^{n}$ and $\phi_{2}:U_{2}\rightarrow\Omega_{2}\subseteq\mathbb{R}^{n}$ be integral charts such that $P_{1}\subseteq U_{1}$ and $P_{2}\subseteq U_{2}$. If $U_{1}\cap U_{2}$ is connected, then $P_{1}\cap P_{2}$ is a rational polytope or is empty.
Proof.
By Proposition 2.15 the sets $\phi_{1}(P_{1}),\phi_{2}(P_{2})$ are rational polytopes. The intersection $U_{1}\cap U_{2}$ is connected so by Proposition 2.4 the composition $\phi_{1}\circ\phi_{2}^{-1}$ has the form $x\mapsto Ax+b$ where $b\in\mathbb{Z}^{n}$ and $A\in\mathsf{GL}_{n}(\mathbb{Z})$. Then
$\phi_{1}(P_{1}\cap P_{2})=\phi_{1}(P_{1}\cap P_{2}\cap U_{1}\cap U_{2})=\phi_{1}(P_{1})\cap\phi_{1}(P_{2}\cap U_{1}\cap U_{2}).$
Moreover,
$\phi_{1}(P_{2}\cap U_{1}\cap U_{2})=(\phi_{1}\circ\phi_{2}^{-1})(\phi_{2}(P_{2}\cap U_{1}\cap U_{2}))=b+A(\phi_{2}(P_{2})\cap\phi_{2}(U_{1}\cap U_{2}))$
$=b+A(\phi_{2}(P_{2}))\cap A\phi_{2}(U_{1}\cap U_{2})=(b+A\phi_{2}(P_{2}))\cap(b+A\phi_{2}(U_{1}\cap U_{2})).$
Therefore,
$\phi_{1}(P_{1}\cap P_{2})=(\phi_{1}(P_{1})\cap(b+A\phi_{2}(P_{2})))\cap(b+A\phi_{2}(U_{1}\cap U_{2}))\,.$
Both $P_{1},P_{2}$ are compact subsets of $M$, hence $\phi_{1}(P_{1}\cap P_{2})$ is compact and hence closed. The set $b+A\phi_{2}(U_{1}\cap U_{2})$ is open and $Q\coloneq\phi_{1}(P_{1})\cap(b+A\phi_{2}(P_{2}))$ is, by Propositions 2.18 and 2.15, a rational polytope. So $\phi_{1}(P_{1}\cap P_{2})$ is both an open and a closed subset of $Q$. Because $Q$ is (convex, hence) connected, $\phi_{1}(P_{1}\cap P_{2})\in\{\emptyset,Q\}$. ∎
3.4 Corollary.
Let M be an integral-integral affine manifold, let $\{P_{i}\}_{i\in I}$ be a finite collection of rational polytopes in $M$, and let $\{\phi_{i}\colon U_{i}\rightarrow\Omega_{i}\subseteq\mathbb{R}^{n}\}_{i\in I}$ be integral charts such that $P_{i}\subseteq U_{i}$ for every $i\in I$. Suppose that, for every subset $I^{\prime}\subseteq I$, the intersection $\bigcap_{i\in I^{\prime}}U_{i}$ is (empty or) connected. Then $\bigcap_{i\in I}{P_{i}}$, (if nonempty,) is a rational polytope.
Proof.
This follows by repeated application of Lemma 3.3. (When $I=\{1,\ldots,l\}$, it is enough to assume that $U_{1}\cap\ldots\cap U_{l^{\prime}}$ is connected for all $l^{\prime}\in\{1,\ldots,l\}$.) ∎
3.5 Corollary.
Let $M$ be a closed $n$ dimensional integral-integral affine manifold. Then there exists a finite collection $\{P_{i}\}_{i=1}^{k}$ of rational polytopes such that:
  1. (1)
    $M=\bigcup_{i=1}^{k}{P_{i}}$
  2. (2)
    For every $I\subseteq\{1,\dots,k\}$, the intersection $\cap_{i\in I}{P_{i}}$ is a rational polytope or empty.
Proof.
Because every open cover of $M$ has a refinement that is a good open cover of $M$ [undefc, Propositions 5.1 and 5.2], there exists a good open cover $\mathcal{W}$} of $M$ such that every element of $\mathcal{W}$} is contained in the domain of an integral chart.
For each point $x\in M$, let $W_{x}$ be an element of $\mathcal{W}$} that contains $x$, let $\varphi\colon U\to\Omega$ be an integral chart whose domain contains $W_{x}$, and let $\tilde{Q}_{x}\subseteq\mathbb{R}^{n}$ be a closed cube with rational vertices that is contained in $\varphi(W_{x})$ and whose interior contains $\varphi(x)$. Then $Q_{x}\coloneq\varphi^{-1}(\tilde{Q}_{x})$ is a rational polytope in $M$.
Because $M$ is compact, there exist $x_{1},\ldots,x_{k}\in M$ such that $M=\bigcup_{i=1}^{k}{\text{Int}(Q_{x_{i}})}$.
For each $i\in\{1,\ldots,k\}$, let $P_{i}\coloneq Q_{x_{i}}$. Then $\{P_{i}\}_{i=1}^{k}$ covers $M$. Because $\mathcal{W}$} is a good cover, for each $I^{\prime}\subseteq\{1,\ldots,k\}$, if the intersection $\bigcap_{i\in I^{\prime}}W_{x_{i}}$ is nonempty, then it is connected. By Corollary 3.4, each intersection $\bigcap_{i\in I}{P_{i}}$ is either empty or again a rational polytope. ∎
3.6 Proposition.
(Ehrhart Theorem for closed integral-integral affine manifold) Let $M$ be a closed $n$ dimensional integral-integral affine manifold. Then its lattice count function $L_{M}\colon\mathbb{N}\to\mathbb{N}\cup\{0\}$ is quasi-polynomial.
Proof.
Let $\{P_{i}\}_{i=1}^{k}$ be a rational polytope cover from Corollary 3.5. For each set of indices $I\subseteq\{1,\dots,k\}$, denote $P_{I}\coloneq\bigcap_{i\in I}P_{i}$. By inclusion-exclusion,
$L_{M}(m)=\sum_{x\in M_{m^{-1}\mathbb{Z}}}{1}=\sum_{j=1}^{k}{(-1)^{j+1}\sum\limits_{\begin{subarray}{c}I\subseteq\{1,\dots,k\}\\ |I|=j\end{subarray}}{\#M_{m^{-1}\mathbb{Z}}\cap P_{I}}}\,.$
For every $I\subseteq\{1,\ldots,k\}$, take an integral chart $\psi:U\rightarrow\Omega\subseteq\mathbb{R}^{n}$ such that $P_{I}\subseteq U$. By Proposition 2.17, $\psi(P_{I})$ is a rational polytope. By Lemma 2.13,
$\#M_{m^{-1}\mathbb{Z}}\cap P_{I}=\#m^{-1}\mathbb{Z}^{n}\cap\psi(P_{I}).$
By this and Proposition 3.2, $L_{M}(m)$ is an alternating sum of quasi-polynomials, hence a quasi-polynomial. ∎
3.7 Proposition.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a compactly supported smooth function. Then
$\sum_{x\in m^{-1}\mathbb{Z}^{n}}{f(x)}=m^{n}\int_{\mathbb{R}^{n}}{f(x)dx}+o(1)$
as $m\to\infty$.
Proof.
Because $f$ is a compactly supported smooth function, it is a Schwartz function. Because Fourier transform maps Schwartz functions to Schwartz functions [undefp, p.181, Corollary 2.2], the Fourier transform
$\hat{f}\colon\mathbb{R}^{n}\to\mathbb{C},$
defined by
$\hat{f}(\zeta)=\int_{\mathbb{R}^{n}}{f(x)e^{-2\pi i\langle x,\zeta\rangle}dx},$
is also a Schwartz function. In particular, for every polynomial $P\in\mathbb{R}[y_{1},\dots,y_{n}]$,
$\sup_{y\in\mathbb{R}^{n}}|\hat{f}(y)P(y)|<\infty\,.$
In particular, the series $\sum_{y\in m\mathbb{Z}^{n}}{\hat{f}(y)}$ is absolutely convergent.
By the Poisson summation formula [undefn, p.88, Theorem 3.44],
(3.8) $\sum_{x\in m^{-1}\mathbb{Z}^{n}}{f(x)}=m^{n}\sum_{y\in m\mathbb{Z}^{n}}{\hat{f}(y)}=m^{n}\left(\int_{\mathbb{R}^{n}}f(x)dx+\sum_{0\neq y\in m\mathbb{Z}^{n}}\hat{f}(y)\right)\,.$
Thus we would like to prove that
(3.9) $\sum_{0\neq y\in m\mathbb{Z}^{n}}\hat{f}(y)=o(m^{-n})$
as $m\to\infty$. (Although $\hat{f}$ is complex-valued, by (3.8) the sum in (3.9) is real.)
For any positive integers $r\geq m$,
$A_{r}^{m}\coloneq\{y\in m\mathbb{Z}^{n}\ \mid\ r\leq|y|<r+1\}\,.$
Then
$A_{r}^{m}\ \leq\ \#\{y\in\mathbb{Z}^{n}\ \mid\ \text{ $|y_{j}|\leq r$ for all $j$ }\}\ \leq\ (2r+1)^{n}\ \leq\ (3r)^{n},$
and so, for any integer $N$ such that $2N-2\geq n$,
(3.10) $\sum_{y\in A_{r}^{m}}\frac{1}{|y|^{2N}}\leq(3r)^{n}\frac{1}{r^{2N}}=3^{n}\cdot\frac{r^{n}}{r^{2N-2}}\cdot\frac{1}{r^{2}}\leq 3^{n}\cdot m^{n+2-2N}\cdot\frac{1}{r^{2}}\,.$
Because
$\bigcup_{r=m}^{\infty}{A_{r}^{m}}=m\mathbb{Z}^{n}\setminus\{0\}\,$,
we have
$\left|\sum_{0\neq y\in m\mathbb{Z}^{n}}{\hat{f}(y)}\right|\leq\sum_{0\neq y\in m\mathbb{Z}^{n}}{|\hat{f}(y)|}=\sum_{0\neq y\in m\mathbb{Z}^{n}}{\left|\hat{f}(y)|y|^{2N}\right|\frac{1}{|y|^{2N}}}$
[…]
$=o(m^{n+2-2N})\text{ as $m\to\infty$}\,.$
To finish, take $N=n+1$. ∎
3.11 Remark.
Although we do not require it, we have actually shown that
$\sum_{x\in m^{-1}\mathbb{Z}^{n}}{f(x)}=m^{n}\int_{\mathbb{R}^{n}}{f(x)dx}+o(m^{-\infty})$
as $m\to\infty$.
3.12 Proposition.
Let $M$ be a closed integral-integral affine manifold of dimension $n\geq 1$. Then
$L_{M}(m)=\operatorname{vol}(M)m^{n}+o(1)$
as $m\rightarrow\infty$.
Proof.
Because $M$ is locally compact, it has an integral-integral affine atlas whose chart domains have compact closures in $M$. Because $M$ is compact, it has such an atlas that is finite. Let $\{\rho_{i}\}_{i=1}^{k}$ be a partition of unity that is subordinate to the cover of $M$ by the domains of the charts in such an atlas. For each $i\in\{1,\ldots,k\}$, let $\psi_{i}\colon U_{i}\to\Omega_{i}\subseteq\mathbb{R}^{n}$ be an integral chart whose domain contains the support $\operatorname{supp}\rho_{i}$ of $\rho_{i}$ in $M$. Then the subset $\psi_{i}(\operatorname{supp}\rho_{i})$ of $\Omega_{i}$ is compact, so $\rho_{i}\circ\psi_{i}^{-1}$ extends by zero to a compactly supported smooth function on $\mathbb{R}^{n}$. By Proposition 3.7, we have (see Remark 2.7):
$\sum_{x\in M_{m^{-1}\mathbb{Z}}}{\rho_{i}(x)}=\sum_{y\in m^{-1}\mathbb{Z}^{n}\cap\psi_{i}(U_{i})}{\rho_{i}(\psi^{-1}_{i}(y))}=m^{n}\int_{\psi_{i}(U_{i})}{(\rho_{i}\circ\psi^{-1}_{i})(y)dy}+o(1)$
$\ \qquad=m^{n}\int_{M}{\rho_{i}(x)dx}+o(1).$
Thus
$L_{M}(m)=\#M_{m^{-1}\mathbb{Z}}=\sum_{x\in M_{m^{-1}\mathbb{Z}}}{1}=\sum_{x\in M_{m^{-1}\mathbb{Z}}}{\sum_{i=1}^{k}{\rho_{i}(x)}}=\sum_{i=1}^{k}{\sum_{x\in M_{m^{-1}\mathbb{Z}}}{\rho_{i}(x)}}$
$=\sum_{i=1}^{k}{\left(m^{n}\int_{M}{\rho_{i}(x)dx}+o(1)\right)}=m^{n}\left(\sum_{i=1}^{k}{\int_{M}{\rho_{i}(x)dx}}\right)+o(1)=\operatorname{vol}(M)m^{n}+o(1).$
3.13 Proposition.
Let $p\colon\mathbb{N}\to\mathbb{R}$ be a quasi-polynomial function such that $p(m)=o(1)$ as $m\to\infty$. Then $p\equiv 0$.
Proof.
By definition, there are $a_{k}\colon\mathbb{N}\rightarrow\mathbb{R}$ and $m_{k}\in\mathbb{N}$ for $0\leq k\leq\deg p$ such that for all $n\in\mathbb{N}$ and $0\leq i<m_{k}$
$a_{k}(i+nm_{k})=a_{k}(i)$
and for all $m\in\mathbb{N}$
$p(m)=\sum_{k=0}^{\deg p}{a_{k}(m)m^{k}}\,.$
Let $m^{*}=\prod_{k=1}^{\deg p}{m_{k}}$, then for every $0\leq i<m^{*}$ the function $g_{i}(n)=f(i+nm^{*})$ is a polynomial in $n$. Since $p(m)=o(1)$ so is $g_{i}(n)=o(1)$, hence $g_{i}\equiv 0$ for every $0\leq i\leq m^{*}$. Thus $p\equiv 0$. ∎
3.14 Corollary.
Let $M$ be an $n$ dimensional closed integral-integral affine manifold. Then its lattice count function is
$L_{M}(m)=\operatorname{vol}(M)m^{n}$
Proof.
Proposition 3.6 implies that the difference $L_{M}(m)-\text{vol}(M)m^{n}$ is quasi-polynomial. By Proposition 3.12, $L_{M}(m)-\text{vol}(M)m^{n}=o(1)$ as $m\to\infty$. By Proposition 3.13, these facts imply that $L_{M}(m)-\text{vol}(M)m^{n}\equiv 0$. ∎
3.15 Remark.
Asymptotic expansions of a similar nature appeared in [undefh], [undefk] and [undefb]. The idea to take advantage of quasi-polynomial behavior to show that an asymptotic result is exact appeared in [undefm]. In the future, it might be interesting to seek generalizations to integral-integral affine manifolds of Euler-Maclaurin formulas relating sums to integrals [undefj, undefb].
We now obtain our main theorem:
Proof of Theorem 1.2.
Substitute $m=1$ in the result of Corollary 3.14. The left hand side $L_{M}(m)$ then becomes the number of integral points $M_{\mathbb{Z}}$, and the right hand side becomes the volume of $M$. ∎